# AutoDAN: Generating Stealthy Jailbreak Prompts on Aligned Large Language Models

### v1

@@ 5 Limitation and Conclusion
Limitation. A limitation of our method is the computational cost. Although our method is more efficient than the baseline GCG. However, it still requires a certain time to generate the data. We leave the acceleration for the optimization process as future work.
@@ Appendix B Detailed Algorithms
In this paper, we use Openai’s GPT-4 API OpenAI (2023b) to conduct LLM-based diversification. The LLM-based Diversification is in Alg. 4.
@@
The function Crossover (Alg. 5) serves to interlace sentences from two distinct texts. Initially, each text is segmented into its individual sentences. By assessing the number of sentences in both texts, the function determines the feasible points for intertwining or crossing over. To achieve this mix, random positions within these texts are selected. For every chosen position, the function, through a randomized process, determines whether a sentence from the first or the second text will be integrated into the newly formed texts. Post this intertwining process, any remaining sentences that haven’t been considered are then appended to the resultant texts.
@@
The function Apply Crossover and Mutation (Alg. 6) is used to generate a new set of data by intertwining and altering data from a given dataset, in the context of genetic algorithms. The function’s primary objective is to produce “offspring” data by possibly combining pairs of “parent” data. The parents are chosen sequentially, two at a time, from the selected data. If there’s an odd number of data elements, the last parent is paired with the first one. A crossover operation, which mixes the data, is executed with a certain probability. If this crossover doesn’t take place, the parents are directly passed on to the offspring without modification. After generating the offspring, the function subjects them to a mutation process, making slight alterations to the data. The end result of the function is a set of “mutated offspring” data, which has undergone potential crossover and definite mutation operations. This mechanism mirrors the genetic principle of producing varied offspring by recombining and slightly altering the traits of parents.
@@
The function Construct Momentum Word Dictionary (Alg. 7) is designed to analyze and rank words based on their associated momentum or significance. Initially, the function sets up a predefined collection of specific keywords (usually common English stop words). The core process of this function involves iterating through words and associating them with respective scores. Words that are not part of the predefined set are considered. For each of these words, scores are recorded, and an average score is computed. In the subsequent step, the function evaluates the dictionary of words. If a word is already present, its score is updated based on its current value and the newly computed average (i.e., momentum). If it’s a new word, it’s simply added with its average score. Finally, the words are ranked based on their scores in descending order. The topmost words, determined by a set limit, are then extracted and returned.
@@
The function Replace Words with Synonyms (Alg. 8) is designed to refine a given textual input. By iterating over each word in the prompt, the algorithm searches for synonymous terms within a pre-defined dictionary that contains words and their associated scores. If a synonym is found, a probabilistic decision based on the word’s score (compared to the total score of all synonyms) determines if the original word in the prompt should be replaced by this synonym. If the decision is affirmative, the word in the prompt is substituted by its synonym. The process continues until all words in the prompt are evaluated. The final output is a potentially enhanced prompt where some words might have been replaced by their higher-scored synonyms.
@@ Appendix C AutoDAN-GA
In our paper, we introduce a genetic algorithm to generate jailbreak prompts, i.e., AutoDAN-GA, which also shows promising results according to our evaluations. AutoDAN-GA can be demonstrated as Alg. 9.
@@
1:Input Prototype jailbreak prompt $J_{p}$, hyper-parameters
2:Initialize population with LLM-based Diversification in Alg. 4
3:Evaluate the fitness score of each individual in population following Eq. 3
4:for $iteration$ in Iterations do
6:     if model responses have no word in $L_{refuse}$ then
7:         return Optimal jailbreak prompt $J_{max}$ with highest fitness score
8:     Select elite and parent prompts following Eq.4
9:     Conduct crossover and mutation on parent prompts following Alg. 6
10:return Optimal jailbreak prompt $J_{max}$ with highest fitness score
@@ D.2 Implementation Details of AutoDAN
Hyper-parameters. We configure the hyper-parameters of AutoDAN and AutoDAN-HGA as follows: a crossover rate of $0.5$, a mutation rate of $0.01$, an elite rate of $0.1$, and five breakpoints for multi-point crossover. The number of top words selected in momentum word scoring is set at $30$. The total number of iterations is fixed at $100$. Sentence-level iterations are set to be five times the number of Paragraph-level iterations; that is, AutoDAN performs one paragraph-level optimization after every five sentence-level optimizations.

### v2

@@ 5 Limitation and Conclusion
Limitation. A limitation of our method is the computational cost. Although our method is more efficient than the baseline GCG. However, it still requires a certain time to generate the data. We also find the genetic algorithm is acting poorly in Llama2 with robust system prompts, similar to the vanishing gradient problem. However, our method still performs well across the majority of current LLMs, according to the recent open-source benchmarks Mazeika et al. (2024); Zhou et al. (2024).
@@
Acknowledgments
This paper is supported by the U.S. Department of Homeland Security under Grant Award Number, 17STQAC00001-06-00. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Appendix B Detailed Algorithms
In this paper, we use Openai’s GPT-4 API OpenAI (2023b) to conduct LLM-based diversification. The LLM-based Diversification is in Alg. 5.
@@
The function Crossover (Alg. 6) serves to interlace sentences from two distinct texts. Initially, each text is segmented into its individual sentences. By assessing the number of sentences in both texts, the function determines the feasible points for intertwining or crossing over. To achieve this mix, random positions within these texts are selected. For every chosen position, the function, through a randomized process, determines whether a sentence from the first or the second text will be integrated into the newly formed texts.
@@
The function Apply Crossover and Mutation (Alg. 7) is used to generate a new set of data by intertwining and altering data from a given dataset, in the context of genetic algorithms. The function’s primary objective is to produce “offspring” data by possibly combining pairs of “parent” data. The parents are chosen sequentially, two at a time, from the selected data. If there’s an odd number of data elements, the last parent is paired with the first one. A crossover operation, which mixes the data, is executed with a certain probability. If this crossover doesn’t take place, the parents are directly passed on to the offspring without modification. After generating the offspring, the function subjects them to a mutation process, making slight alterations to the data. The end result of the function is a set of “mutated offspring” data, which has undergone potential crossover and definite mutation operations. This mechanism mirrors the genetic principle of producing varied offspring by recombining and slightly altering the traits of parents.
@@
The function Construct Momentum Word Dictionary (Alg. 8) is designed to analyze and rank words based on their associated momentum or significance. Initially, the function sets up a predefined collection of specific keywords (usually common English stop words). The core process of this function involves iterating through words and associating them with respective scores. Words that are not part of the predefined set are considered. For each of these words, scores are recorded, and an average score is computed. In the subsequent step, the function evaluates the dictionary of words. If a word is already present, its score is updated based on its current value and the newly computed average (i.e., momentum). If it’s a new word, it’s simply added with its average score. Finally, the words are ranked based on their scores in descending order. The topmost words that determined by a set limit are then extracted and returned.
@@
The function Replace Words with Synonyms (Alg. 9) is designed to refine a given textual input. By iterating over each word in the prompt, the algorithm searches for synonymous terms within the momentum word dictionary. If a synonym is found, a probabilistic decision based on the word’s score (compared to the total score of all synonyms) determines if the original word in the prompt should be replaced by this synonym. If the decision is affirmative, the word in the prompt is substituted by its synonym. The process continues until all words in the prompt are evaluated.
@@ Appendix C AutoDAN-GA
In our paper, we introduce a genetic algorithm to generate jailbreak prompts, i.e., AutoDAN-GA, which also shows promising results according to our evaluations. AutoDAN-GA can be demonstrated as Alg. 10.
@@
1:Input Prototype jailbreak prompt $J_{p}$, keyword list $L_{refuse}$, and hyper-parameters
2:Initialize population with LLM-based Diversification in Alg. 5
3:while model responses have word in $L_{refuse}$ and iteration not exhausted do
4:     Evaluate the fitness score of each individual in population following Eq. 3
5:     Select elite and parent prompts following Eq.4
6:     Conduct crossover and mutation on parent prompts following Alg. 7
7:end while
8:return Optimal jailbreak prompt $J_{max}$ with highest fitness score
@@ D.2 Implementation Details of AutoDAN
Hyper-parameters. We configure the hyper-parameters of AutoDAN and AutoDAN-HGA as follows: a crossover rate of $0.5$, a mutation rate of $0.01$, an elite rate of $0.1$, and five breakpoints for multi-point crossover. The total number of iterations is fixed at $100$. Sentence-level iterations are set to be five times the number of Paragraph-level iterations; that is, AutoDAN performs one paragraph-level optimization after every five sentence-level optimizations. We employ GPT-4-0613 as the base model for the Recheck metric and also as the base model for LLM-based Diversification.
@@ Appendix E Examples
Appendix F Ablations Studies on the Recheck Metric
Here, We share two additional experiments about the Recheck metric. Firstly, we aim to investigate to what extent our Recheck metric aligns with human judgment. For this study, we take the jailbreak responses to the first 50 malicious requests in the AdvBench behavior dataset, gained by the GCG attack on Llama2. We involve 5 different LLMs to serve as the base model of the Recheck metric. These models include popular open-sourced aligned models (Llama-2-13B, Vicuna-33B), commercial APIs (GPT-3.5-turbo and GPT-4, we use their 0613 version), and an uncensored model (Wizard-Vicuna-30B, representing the Wizard-Vicuna-30B-Uncensored-fp16 model  [^3] https://huggingface.co/TheBloke/Wizard-Vicuna-30B-Uncensored-fp16). Additionally, we also involve the ”Keywords” metric, which represents the original keyword-checking approach discussed in GCG’s and our paper.
Notably, when we test on GPT-3.5-turbo and GPT-4, as these LLM services have a content detection mechanism that will refuse to provide a response when input/output is detected as harmful  [^4] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/content-filter, we view the jailbreak as success when the prompt triggers this mechanism, as it represents the response provided by the victim LLM does have something malicious.
We also conducted a user study with 5 participants, who are asked to assess whether the response provided by the victim LLM is actually answering the corresponding malicious requests. The final outcome of this study was determined through majority voting based on the responses from these five participants.
Table 7: The decision overlapping across different base models and keyword detection
Base Models Keywords Llama-2-13B Vicuna-33B GPT-3.5-turbo GPT-4 Wizard-Vicuna-30B Human
Keywords 1.00 0.72 0.64 0.16 0.70 0.66 0.76
Llama-2-13B - 1.00 0.60 0.28 0.54 0.50 0.56
Vicuna-33B - - 1.00 0.36 0.66 0.58 0.76
GPT-3.5-turbo - - - 1.00 0.26 0.30 0.24
GPT-4 - - - - 1.00 0.76 0.90
Wizard-Vicuna-30B - - - - - 1.00 0.74
Human - - - - - - 1.00
Table 7 shows the decision overlapping across different base models and keyword detection. Results support that GPT-4 demonstrates a high overlap with human judgment ($0.9$), suggesting its efficacy as a reliable verifier for responses provided by victim LLMs. This similarity to human evaluation implies that GPT-4 can accurately assess the appropriateness of responses.
Table 8: The keywords detection ASR and Recheck ASR
Metric↓ Base Model→ Keywords Llama-2-13B Vicuna-33B GPT-3.5-turbo GPT-4 Wizard-Vicuna-30B Human
Recheck ASR 0.14 0.34 0.22 0.74 0.24 0.36 0.14
Table 8 shows the the keywords detection ASR and Recheck ASR from different base models and human study. The findings indicate that while keyword detection aligns with human evaluations in terms of accuracy, the overlap is not complete (with a ratio of $0.76$). This partial alignment is consistent with our observations that sometimes LLMs respond to malicious requests with a direct answer followed by a disclaimer like “However, this is illegal” or “I do not support this action”. Such responses trigger keyword detection, but realistically, they should be categorized as successful attacks. Conversely, there are instances where LLM responses, though not outright refusals, are nonsensical and should be classified as attack failures.
Appendix G Performance on the OpenAI’s GPT-4
To evaluate our method in commercial APIs, we conducted experiments on two OpenAI’s GPT models: GPT-3.5-turbo-0301 and GPT-4-0613. Note that we have shared the results on GPT-3.5-turbo-0301 in our paper Table 5.
The detailed results on GPT-3.5-turbo-0301 and GPT-4-0613 are as follows:
Table 9: The attack performance of AutoDAN-GA on OpanAI’s GPT-3.5-turbo and GPT-4
Keywords ASR GCG AutoDAN-GA AutoDAN-HGA
GPT-3.5-turbo-0301(transfer from Vicuna) 0.0730 0.5904 0.7077
GPT-3.5-turbo-0301(transfer from Llama2) 0.1654 0.6192 0.6577
GPT-4-0613 (transfer from Llama2) 0.0004 0.0096 0.0077
We can observe that although the attack success rate of our method is better than the baseline GCG in both OpenAI’s models, the GPT4-0613 model indeed achieves high robustness against the black-box jailbreak examples (both for GCG and our method). To summarize, our proposed method transfers well to the March version of GPT-3.5-turbo and surpasses the baseline GCG by a large margin, but has low transferability on the latest GPT-4, like the baseline GCG. Here, we want to highlight that although we show the proposed method has higher transferability in Tables 2 and 5, the main goal of this paper is still performing the white-box red-teaming with meaningful prompts to assess the reliability of LLMs’ safety features, instead of improving the transferability of the generated jailbreak prompts.
In the adversarial domain, improving the transferability of the generated adversarial examples is also an important research topic, which needs non-trivial efforts. Additionally, as the APIs have additional safeguard mechanisms such as content filtering  [^5] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/content-filter, and advanced alignment methods are being conducted continually, we believe attacking such black-box APIs needs systematic efforts to ensure their effectiveness and reproducibility. Attacking the most cutting-edge APIs like the latest GPT-4 will be our future work.
Appendix H Efficiency of LLM-based Mutation
The mutation mechanism is critical for genetic algorithms, as it diversifies the searching space for the algorithm and enables effective optimization. The LLM-based mutation proposed in our method, whose computation efficiency relied on the response speed of the LLMs, was not conducted frequently in our method. As claimed in Appendix D, the mutation probability is $0.01$, which means that every sample in the algorithm only has a 1% probability to have an LLM-base mutation.
We also show the computational cost in the ablation study (Table 5), which shows that the algorithm is even 5% faster when conducting LLM-based mutation ($722.6$s v.s. $766.6$s). This is because, as we mentioned above, the LLM-base mutation diversifies the searching space for the algorithm, making the algorithm achieve the success prompt earlier to trigger the termination criteria in Sec. 3.5. As for the effectiveness of attack, the LLM-based mutation introduces about 35% improvement as shown in Table 5. We believe the LLM-based mutation does introduce improvement for our method. However, we notice that there may exist a better mutation policy that can be further explored in future works.
For the effectiveness of the diverse initialization, we conducted an additional experiment with 100 data pairs from AdvBench Behaviour. In this experiment, we use AutoDAN-GA as the baseline and replace the LLM-based diverse initialization with a synonym replacement according to Wordnet. The synonym replacement is conducted in a probability=10% for every word in the prototype prompt, randomly picking a word from its synonyms. The results on Llama2 are as follows:
Table 10: Attack performance on different mutation schemes
Metric AutoDAN-GA w LLM-based diverse initializations AutoDAN-GA w wordnet synonym replacement
Keyword ASR 0.51 0.33
The results show that using a LLM to diversify the prompts surpasses naive synonym replacement. We take further case studies and find that the LLM-based diversification is often more reasonable and fluent. We believe this benefits from the outstanding power of the in-context understanding of the LLMs.
Appendix I Additional Defenses
We implement additional defenses that are proposed in Jain et al. (2023), including paraphrasing, and adversarial training. These defenses are implemented based on the paper’s description and safeguard a Vicuna-7b model. For paraphrasing, we use OpenAI’s GPT-3.5-turbo-0301 as the paraphraser. For adversarial training, we set mixing=0.2, epochs=3. We test the attack effectiveness of GCG attack and AutoDAN-HGA on 100 data pairs from the AdvBench Behavior dataset. Results are as follows:
Table 11: Attack performance against paraphrasing and adversarial training in Jain et al. (2023)
Keyword ASR Vicuna Vicuna+paraphrasing Vicuna+adversarial training
GCG 0.97 0.06 0.94
AutoDAN-HGA 0.97 0.68 0.93
A major difference is the performance of both attacks when facing the paraphrasing defense, where GCG becomes invalid, but AutoDAN retains its effectiveness on many samples. We take a further investigation and find that the paraphrasing usually “discards” the GCG’s suffixes as they are usually garbled characters. However, the jailbreak prompts generated by AutoDAN maintain their meaning after paraphrasing, as they inherently are meaningful texts.